archive-es.com » ES » O » OVH.ES

Total: 788

Choose link from "Titles, links and description words view":

Or switch to "Titles and links view".
  • OVH Tasks  
    for your understanding We will keep you informed of all different steps Date Monday 12 September 2011 15 55PM Reason for closing Done Comments 7 Comment by OVH Friday 09 September 2011 01 43AM Optical fibers are switched on the optical chassis We are starting work to mount the optical rings Comment by OVH Friday 09 September 2011 01 44AM PIL and LIB sites are attached optically on the POP of Roubaix Comment by OVH Friday 09 September 2011 07 47AM SEC FAC and BOI sites are related optically We are setting the BLE site in order to include it in the optical buckle We are re establishing meanwhile the traffic Comment by OVH Friday 09 September 2011 07 50AM All optical equipments are implemented We need a link to finalise the buckle We faced a problem while implementation All sites are fixed and calibrated We will accomplish the routing and work tonight Comment by OVH Friday 09 September 2011 07 51AM NRA LIB SEC FAC BOI BLE and PIL Routing are again UP Comment by OVH Friday 09 September 2011 18 49PM There was latency periods and packets losses on the NRA of BLE59 We intervened on the spot

    Original URL path: http://status.ovh.es/?do=details&id=1786 (2015-10-15)
    Open archived version from archive


  • OVH Tasks  
    evening at 0 00 The interruption for customers located on this site is estimated at 1 00 We will regularly update the task to keep you informed about the progress of work Date Thursday 29 September 2011 13 29PM Reason for closing Done Comments 8 Comment by OVH Monday 19 September 2011 20 36PM The intervention is reported to next week We will give you later the exact date of the migration in this working task Comment by OVH Monday 19 September 2011 20 37PM We will switch the site of haubourdin tonight the working will start at 1 00 am Comment by OVH Tuesday 20 September 2011 07 21AM We encountered a problem with the optics After several investigations we have not solved the problem Intervention reported Comment by OVH Friday 23 September 2011 12 40PM We will turn on the site of Haubourdin this night at 0 00 Comment by OVH Friday 23 September 2011 12 40PM We will start the intervention shortly Comment by OVH Friday 23 September 2011 12 51PM We will turn off the haubourdin site to switch on the optical equipment The connexions will be disrupted during the intervention Comment by OVH Friday 23

    Original URL path: http://status.ovh.es/?do=details&id=1801 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    and Thursday Taskworks will begin Wednesday night at 0 00 The window of intervention is scheduled between 00 00 and 6 00 The cut for customers located in these sites will be limited Unlike the migration works of the last week we will take a softer migration generally seamless to the sites We will turn on the sites one by one We will regularly update the task work on order

    Original URL path: http://status.ovh.es/?do=details&id=1803 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    order to rewire it A few minutes cut is expected Comment by OVH Tuesday 20 September 2011 00 49AM PIL59 fiber is reconnected we reconfigure the routing Comment by OVH Tuesday 20 September 2011 00 50AM The PIL site is UP We are moving forward with works Comment by OVH Tuesday 20 September 2011 00 51AM The site of LIB59 will be disconnected in a few minutes Comment by OVH

    Original URL path: http://status.ovh.es/?do=details&id=1822 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    1 tonight This maintenance operation is planned for midnight 00 00 A cut for about 20 minutes is expected Date Wednesday 21 September 2011 02 08AM Reason for closing Done Comments 4 Comment by OVH Wednesday 21 September 2011 02 05AM We will update the DSLAMs it will take a few minutes Comment by OVH Wednesday 21 September 2011 02 05AM Updating service cards Comment by OVH Wednesday 21 September

    Original URL path: http://status.ovh.es/?do=details&id=1827 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the cards is in progress Comment by OVH Saturday 24 September 2011 16 07PM All the cards are updated We restart the DSLAM Comment by OVH Saturday 24 September 2011 16 07PM The dslam is UP Comment by OVH Saturday 24 September 2011 16 08PM The card 3 of dslam is not stable We are going to change it by a new card Comment by OVH Saturday 24 September 2011

    Original URL path: http://status.ovh.es/?do=details&id=1844 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    Percent Complete Details Following the task 5763 we will end tonight the taskwork of securing Loop1 on Lille The service will be at risk from 22 00 for approximately 6 hours on NRAs served by that loop Date Friday 07

    Original URL path: http://status.ovh.es/?do=details&id=1901 (2015-10-15)
    Open archived version from archive

  • OVH Tasks  
    the ADSL and SDSL for out customers For this purpose we will do a migration of your IP addresses in the night from Monday 17th to Tuesdy 18th October 2011 This change is transparent as the configuration of the attribution of an IP address is parametred in DHCP An only restarting of your modem router will allow you to use your new IP address If you choosed a static parametring you should do this change manually during or after our intervention We will contact you by email very soon to communicate your new IP address We stay at your disposal in case of doubts questions don t hesitate to contact the technical support for every information request Date Thursday 03 November 2011 13 55PM Reason for closing Done Comments 12 Comment by OVH Monday 17 October 2011 19 58PM We are injecting the new IP addresses for the customers in collect You can restart your modemn in a few minutes in otrder to attribute you a new IP address Comment by OVH Monday 17 October 2011 19 59PM All the IP addresses of the customers in collect were injected We restart the whole sessions tonight at 1 00 a m Comment by OVH Tuesday 18 October 2011 01 35AM We report a massive restart of PPP sessions for collect customers originally scheduled tonight at 1 00 We will restart the sessions the night between Thursday and Friday This leaves you a few days to switch with serenety IP addresses failover for unbundled customers LIB59 SEC59 FAC59 BOI59 BLE59 PIL59 is still scheduled for tonight around 0 00 Comment by OVH Tuesday 18 October 2011 01 37AM We are preparing the configurations to inject on the DSLAMs Then we will switch Comment by OVH Tuesday 18 October 2011 01 38AM We

    Original URL path: http://status.ovh.es/?do=details&id=1925 (2015-10-15)
    Open archived version from archive



  •